OpenStudy (anonymous):

Which statement is true for the expression a + (3 + 6y)? It can be written as (a + 3) + 6y by using associative property of addition. It can be written as a + 3 + 6y + (3 + 6y) by using distributive property of addition. It can be written as a + 3 + 6y + (3 + 6y) by using associative property of addition. It can be written as (a + 3) + 6y by using distributive property of addition

mathslover (mathslover):

Associative property of addition states that : \[\huge{(a+b)+c=a+(b+C)}\] hence \[\huge{a+(3+6y)=(a+3)+6y}\] Hence 1st one is right :)

OpenStudy (anonymous):

yay! thats what i guessed thanks!

OpenStudy (anonymous):

Distributive property only applies to multiplication.

mathslover (mathslover):

congrats .. just be confident to your answer and u will get it right .. congo . :)